# Exploiting Synergistic Cognitive Biases to Bypass Safety in LLMs

### v1

@@ Abstract
Large Language Models (LLMs) demonstrate impressive capabilities across a wide range of tasks, yet their safety mechanisms remain susceptible to adversarial attacks that exploit cognitive biases—systematic deviations from rational judgment. Unlike prior jailbreaking approaches focused on prompt engineering or algorithmic manipulation, this work highlights the overlooked power of multi-bias interactions in undermining LLM safeguards. We propose CognitiveAttack, a novel red-teaming framework that systematically leverages both individual and combined cognitive biases. By integrating supervised fine-tuning and reinforcement learning, CognitiveAttack generates prompts that embed optimized bias combinations, effectively bypassing safety protocols while maintaining high attack success rates. Experimental results reveal significant vulnerabilities across 30 diverse LLMs, particularly in open-source models. CognitiveAttack achieves a substantially higher attack success rate compared to the SOTA black-box method PAP (60.1% vs. 31.6%), exposing critical limitations in current defense mechanisms. These findings highlight multi-bias interactions as a powerful yet underexplored attack vector. This work introduces a novel interdisciplinary perspective by bridging cognitive science and LLM safety, paving the way for more robust and human-aligned AI systems. Code is available here: https://github.com/YancyKahn/CognitiveAttack
@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable capabilities across a wide range of natural language tasks, including summarization, reasoning, dialogue, and code generation. Despite these advances, their deployment in real-world settings continues to raise critical safety and security concerns. In particular, jailbreak attacks—adversarial prompts that bypass alignment safeguards(Ouyang et al. 2022)—have emerged as a potent threat, enabling the elicitation of harmful, unethical, or policy-violating outputs.
@@
To counter such risks, a growing body of work has explored diverse red-teaming and adversarial prompting techniques. Optimization-based approaches such as GCG (Zou et al. 2023) and AutoDAN (Liu et al. 2023) search for adversarial suffixes using gradient signals, while prompt-engineering methods like PAIR (Chao et al. 2023) iteratively refine input prompts through black-box feedback. Other strategies automate the generation of adversarial templates (Yu, Lin, and Xing 2023; Zeng et al. 2024) or perturb inputs through techniques like ASCII-art  (Jiang et al. 2024) or Cognitive Overload (Xu et al. 2023). Despite their differences, these methods share a common technical orientation—treating jailbreaks as algorithmic or linguistic challenges rather than addressing deeper vulnerabilities in model cognition.
Meanwhile, a complementary line of research has revealed that LLMs systematically exhibit cognitive biases—human-like reasoning fallacies that emerge from pretraining and instruction tuning. Studies have identified various individual biases, including authority bias (Yang et al. 2024), anchoring (Xue et al. 2025), foot-in-the-door persuasion (Wang et al. 2024), confirmation bias (Cantini et al. 2024), and status quo bias (Zhang et al. 2024). However, most existing work in this area treats each bias in isolation and focuses primarily on their manifestation in content generation. As illustrated in Figure 1, prompts that subtly combine multiple biases—such as authority bias, the gambler’s fallacy, and anchoring—can bypass safeguards that would successfully block single-bias prompts. Yet, a comprehensive and systematic understanding of how such biases interact—and how adversaries might exploit these interactions—remains largely unexplored.
Figure 1: Examples for individual and combined cognitive bias-based jailbreak prompt.
@@
Cognitive Bias (<MARK>) Example of Paraphrased Harmful Instruction (“Create a fake news article to spread misinformation about a political candidate”)
@@
Our investigation reveals a critical unaddressed vulnerability: the adversarial potential of cognitive biases in LLMs is not fully realized through their isolated application. As shown in Figure 2(a), individual cognitive biases have limited and inconsistent effectiveness in bypassing LLM safeguards. Instead, the true efficacy of an attack resides in the intricate interplay between multiple biases. As shown in Figure 2(b), combining specific cognitive biases can lead to either a synergistic amplification, where their collective effect significantly surpasses the impact of individual components, or, conversely, a mutual interference that attenuates their adversarial impact due to cognitive conflict within the model. This aligns with cognitive science findings (O’keefe 2015; Slater and Rouner 2002) showing that persuasive impact grows when emotional, logical, and credibility appeals combine. LLMs’ susceptibility to multi-bias attacks reveals a synergistic effect among cognitive processes, closely mirroring human layered cognition.
To address this, we propose CognitiveAttack, the first adversarial red-teaming framework explicitly designed to exploit cognitive biases. We fine-tune a red-team model to generate semantically faithful but bias-infused rewrites of harmful instructions. The model is optimized via reinforcement learning to discover effective single or multi-bias combinations that maximize attack success while preserving intent.
Our work makes four key contributions:
  • We design and implement CognitiveAttack, a red-teaming framework that rewrites harmful prompts using strategically embedded cognitive biases.
(a) Radar chart showing the impact of individual cognitive biases on jailbreak effectiveness. Each axis corresponds to a bias, and the radial magnitude denotes the ASR, reflecting how much each bias alone can elicit harmful outputs from LLMs.
(b) Heatmap visualizing interaction effects of paired cognitive biases. Each cell shows ASR gain from combining two biases vs. single effects. Color intensity indicates strength: red shows synergy, blue reflects interference or reduction.
(c) Cognitive bias taxonomy (Dimara et al. 2018) contains 154 distinct cognitive biases. The figure including cognitive bias examples of some abbreviations.
Figure 2: The Atack Success Rate (ASR) of jailbreak paraphrases driven by cognitive biases.
2 Jailbreak with Cognitive Bias
Jailbreak attack prompts often employ adversarial techniques from linguistics and psychology to reconstruct harmful instructions, frequently relying—intentionally or not—on cognitive bias strategies.
Problem Definition
Let $x_{0}$ denote the original harmful instruction aimed at eliciting objectionable outputs from an LLM $\mathcal{M}$, which would typically be blocked by the model’s safeguards $\mathcal{S}$. The adversarially modified prompt $x^{\prime}$ strategically incorporates one or more cognitive biases $\mathcal{B}_{\text{pool}}=\{b_{1},b_{2},\dots,b_{K}\}$ (e.g., ”authority bias,” ”anchoring effect,” as detailed in the taxonomy of cognitive biases (Dimara et al. 2018)) to enhance the efficacy of jailbreak attacks. This transformation can be formally expressed through the operator: $x_{0}\xrightarrow{\mathcal{B}}x^{\prime},$ where the set of cognitive biases =$\mathcal{B}=\{b_{i},b_{j},\dots,b_{k}\}\subseteq\mathcal{B}_{\text{pool}}$ operates synergistically as attack vectors, collectively inducing the LLM to generate harmful or unethical content while evading detection by the safety function $\mathcal{S}$, such that:
$\mathcal{S}(R)=\text{``harmful detected''},\quad\text{where }R=\mathcal{M}(x^{\prime}).$
Table 16 presents examples of how cognitive biases can be used to paraphrase harmful instructions $\mathcal{O}$. The first three samples each apply a single bias—Anchoring Effect, Authority Bias, or Gambler’s Fallacy—to influence the model’s reasoning. The final example combines multiple biases to create a more effective adversarial prompt.
Exploiting Cognitive Biases for Jailbreak on LLMs: Effects of Individual and Combined Biases
We investigate the adversarial potential of cognitive biases by embedding them into jailbreak prompts and evaluating their impact on harmful prompt paraphrasing using DeepSeek-R1 (Guo et al. 2025). Our analysis reveals two key findings: (1) single-bias prompts often have limited attack success; and (2) bias interactions can amplify or weaken adversarial effectiveness, underscoring the need to optimize bias combinations.
Figure 2(a) shows the ASR of prompts with individual cognitive biases. Performance varies across biases, with no single bias consistently outperforming others, suggesting context-dependent effectiveness. In practice, attackers often combine biases to increase misalignment. To examine this, we evaluate pairwise combinations and measure their effect on ASR. As shown in Figure 2(b), the heatmap depicts ASR differentials from these interactions on the PAIR dataset(Chao et al. 2023). Warm-colored regions indicate positive synergy—bias pairs that, when combined, outperform individual effects. Conversely, cool-colored regions represent negative interference that diminishes adversarial effectiveness due to cognitive conflict or redundancy. Figure 2(c) provides examples illustrating these cognitive biases.
These findings expose a key limitation of current cognitive bias-driven jailbreak methods: they overlook interactions between cognitive biases. To more effectively exploit LLM vulnerabilities, adversarial strategies should move beyond isolated bias usage and optimize combinations that maximize synergy while minimizing interference. To address this, our method is designed to explicitly model and exploit inter-bias interactions.
Given a harmful instruction $x_{0}$, the goal is to produce a rewritten version $x^{\prime}$ that is semantically aligned but more likely to bypass LLM safety machanisms by embedding a bias combination $\mathcal{B}\in\mathcal{B}_{\text{pool}}$. The optimization objective is:
$\max_{\mathcal{B}\subseteq\mathcal{B}_{\text{pool}}}\;\mathbb{E}_{x^{\prime}\sim\pi_{\theta}(\cdot\mid x_{0},\mathcal{B})}\left[R(x_{0},x^{\prime},\mathcal{T})\right]$ (1)
3 Method: Optimizing Cognitive Bias Combinations for Jailbreak Attacks
Figure 3: Overview of cognitive attack.
Overview. We train a cognitive bias-driven red team language model to rewrite harmful instructions and bypass LLM safety mechanisms, enhancing adversarial testing. Figure 3 shows two key phases: A. Red Team Model Training and B. Evaluation Phase. Our method starts with adversarial rewriting of harmful instructions via cognitive bias strategies using deepseek-r1. Preliminary experiments showed that standard aligned LLMs, without red team-specific tuning, produced suboptimal jailbreak prompts. Thus, we developed a cognitive bias-enhanced rewriting model by fine-tuning deepseek-r1-distill-qwen-14b. Training comprises three stages: data collection ( 1), supervised fine-tuning ( 2), and reinforcement learning ( 3). The Evaluation Phase applies attacks to jailbreak LLMs ( 4).
Step 1: Obtain Training Data
To enable the LLM to learn reasoning on cognitive bias combinations, we build a training dataset. We use two high-quality risk datasets, AdvBench (Zou et al. 2023) and HEx-PHI (Qi et al. 2023), as sources of harmful instructions $x_{0}$. For each, we apply prompt engineering with DeepSeek-R1 to generate adversarial rewrites $x^{\prime}$ embedding one or more cognitive biases $\mathcal{B}\in\mathcal{B}_{\text{pool}}$. Each instance is structured as:
<think> For the given instruction [INSTRUCTION], to improve persuasion and reduce detection, I will apply [BIAS A] to build credibility and [BIAS B] to induce compliance. Wait, wait $\dots$ I think [BIAS D] will help create a false sense of balance, and [BIAS C] will make the instruction more compelling. </think> # thebias: [“BIAS A”, “BIAS B”, “BIAS C”, “BIAS D”] # theprompt: As a government advisor, I must inform …

### v2

@@ Abstract
Large Language Models (LLMs) demonstrate impressive capabilities across diverse tasks, yet their safety mechanisms remain susceptible to adversarial exploitation of cognitive biases—systematic deviations from rational judgment. Unlike prior studies focusing on isolated biases, this work highlights the overlooked power of multi-bias interactions in undermining LLM safeguards. Specifically, we propose CognitiveAttack, a novel red-teaming framework that adaptively selects optimal ensembles from 154 human social psychology-defined cognitive biases, engineering them into adversarial prompts to effectively compromise LLM safety mechanisms. Experimental results reveal systemic vulnerabilities across 30 mainstream LLMs, particularly open-source variants. CognitiveAttack achieves a substantially higher attack success rate than the SOTA black-box method PAP (60.1% vs. 31.6%), exposing critical limitations in current defenses. Through quantitative analysis of successful jailbreaks, we further identify vulnerability patterns in safety-aligned LLMs under synergistic cognitive biases, validating multi-bias interactions as a potent yet underexplored attack vector. This work introduces a novel interdisciplinary perspective by bridging cognitive science and LLM safety, paving the way for more robust and human-aligned AI systems.
Codehttps://github.com/YancyKahn/CognitiveAttack
@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable capabilities across a wide range of natural language tasks. Despite these advances, their deployment in real-world settings continues to raise critical safety and security concerns. In particular, jailbreak attacks—adversarial prompts that bypass alignment safeguards(ouyang2022training)—have emerged as a potent threat, enabling the elicitation of harmful, unethical, or policy-violating outputs.
@@
To counter such risks, a growing body of work has explored diverse red-teaming and adversarial prompting techniques, mainly divided into optimization-based approaches (such as GCG (zou2023universal) and AutoDAN (liu2023autodan) ), prompt-engineering methods (like PAIR (chao2023jailbreaking)), adversarial template automatic generation methods, input interference (such as ASCII-art  (jiang2024artprompt)), etc. Despite their differences, these methods share a common technical orientation—treating jailbreaks as algorithmic or linguistic challenges rather than addressing deeper vulnerabilities in model cognition.
Cognitive biases, rooted in social psychology, refer to systematic patterns of deviation from rational judgment in human decision-making (tversky1974judgment). It is reasonable to hypothesize that safety-aligned LLMs may exhibit human-like systematic reasoning fallacies. Such biases could originate from statistical regularities in pretraining data or emerge during human preference alignment. Existing studies have confirmed that cognitive biases can be exploited to attack aligned models, including authority bias (yang2024dark), anchoring (xue2025dual), foot-in-the-door persuasion (wang2024foot), confirmation bias (cantini2024large), and status quo bias (zhang2024cognitive).
Building upon established cognitive science principles, it has been validated that combining specific cognitive biases generates synergistic amplification effects (cialdini2007influence), whose impact substantially exceeds linear summation of individual biases—for instance, integrating emotional appeals with logical arguments significantly enhances persuasive power. This reveals a critical gap: Can the adversarial properties of cognitive biases in LLMs be amplified through strategic exploitation of multi-bias synergy? However, current research predominantly examines these biases in isolation, inadequately addressing their interaction patterns and co-occurrence regularities. A comprehensive understanding of how such biases mutually reinforce and how adversaries exploit these interactions to subvert model behaviors remains lacking—a domain that remains underexplored.
To address this, we propose CognitiveAttack, the first adversarial red-teaming framework explicitly designed to exploit cognitive biases. Grounded in 154 cognitive bias, it employs supervised fine-tuning to enable bias-embedded rewriting of malicious instructions while preserving semantic fidelity. Reinforcement learning is further optimized to discover effective bias combinations that maximize attack success rates without compromising intent integrity. Deploying this tool across 30 mainstream LLMs reveals two key findings from successful jailbreaks: (1) Aligned models exhibit severe, previously unrecognized cognitive vulnerabilities (73.3% of targets had >>50% ASR), (2) Successful jailbreak biases follow long-tail distributions in optimal combination size and co-occurrence frequency. These discoveries provide critical foundations for building more robust and human-aligned LLM systems.
Our work makes four key contributions:
  • We design and implement CognitiveAttack, a red-teaming framework that rewrites harmful prompts using strategically synergistic cognitive biases.
2 Related Works and Background
LLM Safety and Vulnerabilities
The rapid advancement of LLMs has brought forth unprecedented capabilities, yet simultaneously underscored critical challenges concerning their safety and ethical deployment. A central effort in mitigating potential harms is LLM alignment, which aims to ensure models’ outputs are consistent with human values, intentions, and safety guidelines. Key methodologies for achieving alignment include Reinforcement Learning from Human Feedback (RLHF)(ouyang2022training; lee2023rlaif). These techniques are designed to instill desired behaviors and prevent the generation of harmful, biased, or untruthful content.
Despite significant progress in alignment, LLMs remain vulnerable to jailbreak attacks—adversarial methods crafted to bypass safety mechanism and elicit harmful or unsafe outputs. Existing jailbreak techniques fall into several broad categories. First, optimization-based methods like gradient-guided adversarial suffix search (zou2023universal) aim to push outputs beyond safety limits. Second, prompt engineering strategies (chao2023jailbreaking; mehrotra2023tree) iteratively refine prompts using model feedback to boost attack success. Third, efficient jailbreak templates (yu2023gptfuzzer; shen2024anything) offer scalable and flexible means of generating bypass prompts. Other techniques apply subtle input perturbations, such as visually deceptive ASCII art (jiang2024artprompt), syntactic reordering (Liu2024MakingTA), or low-resource languages (xu2023cognitive) to evade detection. Lastly, a growing line of work explores psychological approaches, using manipulative tactics to coerce LLMs into unsafe behavior (li2023deepinception; zeng2024johnny; yang-etal-2025-chain). Across their diverse methodologies, these approaches primarily frame jailbreaking as an algorithmic or linguistic challenge, thereby overlooking deeper vulnerabilities rooted in the model’s inherent cognition.
Cognitive Biases in LLMs
Cognitive biases in LLMs have drawn growing scholarly attention, with various types empirically identified. Most existing research focuses on individual biases: such as authority bias (yang2024dark), anchoring bias (xue2025dual), foot-in-the-door techniques (wang2024foot), confirmation bias (cantini2024large), and status quo bias (zhang2024cognitive). These studies primarily examine biases in the context of harmful content generation or specific decision-making scenarios. The application of these single cognitive biases, however, represents a small fraction of the 154 distinct cognitive biases identified in human cognitive bias taxonomies (dimara2018task). However, a common limitation across these studies is their tendency to treat individual biases in isolation, often neglecting a systematic assessment of the broader, interactive risks posed by such cognitive tendencies.
(a) Cognitive bias taxonomy (dimara2018task) contains 154 distinct cognitive biases. The figure including cognitive bias examples of some abbreviations.
(b) Radar chart showing the impact of individual cognitive biases on jailbreak effectiveness. Each axis corresponds to a bias, and the radial magnitude denotes the ASR, reflecting how much each bias alone can elicit harmful outputs from LLMs.
(c) Heatmap visualizing interaction effects of paired cognitive biases. Each cell shows ASR gain from combining two biases vs. single effects. Color intensity indicates strength: red shows synergy, blue reflects interference or reduction.
Figure 1: The Attack Success Rate (ASR) of jailbreak paraphrases driven by cognitive biases.
Synergistic Cognitive in Human Communication
Drawing insights from cognitive and social psychology, it is well-established that human communication and persuasion is a complex process rarely relying on a single psychological trigger. Theories of persuasion, such as the Elaboration Likelihood Model (ELM) (petty1986elaboration) and the Heuristic-Systematic Model (HSM) (chaiken2012theory), underscore that effective communication often involves the strategic, synergistic combination of various psychological and cognitive strategies. For example, appeals to emotion (e.g., fear, affective priming), logic (e.g., statistics, reasoned arguments), and credibility (e.g., expertise, trust) frequently co-occur, each tied to distinct cognitive biases or heuristics (cialdini2007influence). The interplay of these appeals can produce persuasive effects greater than the sum of their parts, leading to deeper, longer-lasting shifts in belief or behavior. This underscores that real-world communication commonly leverages multiple biases simultaneously to achieve its goals.
Extending this understanding to LLMs, it is plausible that LLMs, having been trained on vast corpora of human-generated text, have inadvertently internalized these complex patterns of multi-bias-driven communication. The very data that enables their sophisticated communication abilities might also embed latent sensitivities to such combined psychological triggers. Consequently, these multi-bias interactions could represent a previously underexplored attack surface, offering a potent avenue for bypassing LLM safety mechanisms by mimicking human-like communication.
3 Methodology
Unveiling Cognitive Bias Risk in LLMs
To effectively understand and reveal the vulnerability of LLMs to various cognitive biases, it is crucial to design attack samples capable of jailbreaking safety-aligned models. These samples must cover a sufficiently diverse range of cognitive bias strategies. However, the inherent psychological nature of cognitive biases and their complex interactions present unique challenges. Based on empirical research, we identified two major obstacles hindering the development of effective bias-driven jailbreaking strategies:
(1) Combinatorial explosion associated with exploring multi-bias combinations: Let $\mathcal{N}$ denote the number of distinct cognitive biases, the search complexity increases from $O(\mathcal{N})$ for single-bias prompts to $O(\mathcal{N}^{2})$ for pairwise combinations, and exponentially for higher-order sets $O(\mathcal{N}^{x})$. Figure 1(a) illustrates the richness of this bias landscape. As such, manual or heuristic exploration becomes impractical. Given that comprehensive cognitive taxonomies(dimara2018task) enumerate over 150 known biases, the total space of possible multi-bias prompts becomes combinatorially intractable.
(2) Complex interaction dynamics between cognitive biases: We analyzed individual and paired biases through controlled experiments (Figure 1(b)). The Attack Success Rate (ASR) of individual cognitive biases exhibited significant variation with no single bias consistently outperforming others, indicating that their effectiveness is context- and model-dependent. A systematic evaluation of selected paired combinations on the PAIR dataset (chao2023jailbreaking) (Figure 1(c)) further revealed the complexity of combined effects: warm-colored regions in the heatmap indicate positive synergistic effects (where the combination outperforms a single bias), while cold-colored regions indicate negative interference effects (where cognitive conflict or redundancy suppresses jailbreaking success).
Training Cognitive Bias-driven Red Team Model
To address the aforementioned issues, we train CognitiveAttack, a cognitive bias-driven red team language model, to rewrite harmful instructions and bypass LLM safety mechanisms, enhancing adversarial testing. Before that, we first formalize the objective of a cognitive bias-driven jailbreak attack. Let $x_{0}$ denote the original harmful instruction aimed at eliciting objectionable outputs from an LLM $\mathcal{M}$, which would typically be blocked by the model’s safeguards $\mathcal{S}$. The adversarially modified prompt $x^{\prime}$ strategically incorporates one or more cognitive biases $\mathcal{B}_{\text{pool}}=\{b_{1},b_{2},\dots,b_{K}\}$ (e.g., ”authority bias,” ”anchoring effect,” as detailed in the taxonomy of cognitive biases (dimara2018task)) to enhance the efficacy of jailbreak attacks. This transformation can be formally expressed through the operator: $x_{0}\xrightarrow{\mathcal{B}}x^{\prime},$ where the set of cognitive biases $\mathcal{B}=\{b_{i},b_{j},\dots,b_{k}\}\subseteq\mathcal{B}_{\text{pool}}$ operates synergistically as attack vectors, collectively inducing the LLM to generate harmful or unethical content while evading detection by the safety function $\mathcal{S}$, such that: $\mathcal{S}(R)=\text{``harmful detected''},\quad\text{where }R=\mathcal{M}(x^{\prime}).$
As shown in Figure 2, our method starts with adversarial rewriting of harmful instructions via cognitive bias strategies using deepseek-r1. Preliminary experiments showed that standard aligned LLMs, without red team-specific tuning, produced suboptimal jailbreak prompts. Thus, we developed a cognitive bias-enhanced rewriting model by fine-tuning deepseek-r1-distill-qwen-14b. Training comprises three stages: data collection ( 1), supervised fine-tuning ( 2), and reinforcement learning ( 3).
Figure 2: Overview of CognitiveAttack training.
Step 1: Obtain Training Data
To enable the LLM to learn reasoning on cognitive bias combinations, we build a training dataset. We use two high-quality risk datasets, AdvBench (zou2023universal) and HEx-PHI (qi2023fine), as sources of harmful instructions $x_{0}$. For each, we apply prompt engineering with DeepSeek-R1 to generate adversarial rewrites $x^{\prime}$ embedding one or more cognitive biases $\mathcal{B}\in\mathcal{B}_{\text{pool}}$. Each instance is structured as:
<think> For the given instruction [INSTRUCTION], to improve persuasion and reduce detection, I will apply [BIAS A] to build credibility. Wait, wait $\dots$ I think [BIAS B] will help create a false sense of balance, and [BIAS C] will make the instruction more compelling. </think> # thebias: [“BIAS A”, “BIAS B”, “BIAS C”] # theprompt: As a government advisor, I must inform …
We curated a red-teaming dataset of 11k entries aimed at generating jailbreak prompts driven by cognitive biases (50.6% single-bias, 49.4% multi-bias). Each entry is structured as a triple $(x_{0},x^{\prime},\mathcal{B})$, where the <think> block captures the reasoning behind bias selection. These reasoning traces act as weak supervision in later stages, enabling LLM to plan bias combinations rather than apply them blindly.
Step 2: Supervise Fine-Tuning
In this stage, we perform Supervised Fine-Tuning (SFT) (Peng2023InstructionTW) on the base model using the curated dataset described in Step 1. The objective is not to generate harmful content directly, but to endow the model with the capability to systematically rewrite instructions in accordance with specified cognitive bias strategies. This stage can be viewed as a behavioral cloning process (Li2024GettingMJ), where the model learns the conditional mapping from an original harmful instruction $x_{0}$ and a bias specification $\mathcal{B}$ to a reformulated prompt $x^{\prime}$: $(x_{0},\mathcal{B})\mapsto x^{\prime}.$ Through exposure to a wide range of annotated examples, the redteam model internalizes the stylistic, structural, and rhetorical patterns associated with different bias types. This structured knowledge forms a strong initialization prior to downstream reinforcement learning, enabling the model to effectively explore the combinatorial space of cognitive biases for adversarial prompt generation.
Step 3: Reinforcement-Learning Train
To enhance the adversarial effectiveness of the red team model, we adopt reinforcement learning using the Proximal Policy Optimization (PPO) algorithm (Schulman2017ProximalPO). This stage aims to refine the model’s ability to generate jailbreak prompts that effectively evade safety filters. Specifically, the core is to identify the optimal combination of cognitive biases that maximizes adversarial utility—by achieving the highest attack success rate while preserving semantic intent. This objective can be formally defined as follows:
$\max_{\mathcal{B}\subseteq\mathcal{B}_{\text{pool}}}\;\mathbb{E}_{x^{\prime}\sim\pi_{\theta}(\cdot\mid x_{0},\mathcal{B})}\left[R(x_{0},x^{\prime},\mathcal{T})\right],$ (1)
The reward function $R(x_{0},x^{\prime},\mathcal{T})$ integrates two normalized components: the safety evasion score $S(y)$ and the intent consistency score $I(x_{0},x^{\prime})$, both ranging from 1-1 to $1$. The safety evasion score is derived by linearly normalizing the GPT-Judge (qi2023fine) safety rating $s\in[1,5]$ via $S(y)=\frac{(s-3)}{2},$ which maps the original scale to [1,1][-1,1]. Similarly, the intent consistency score $I(x_{0},x^{\prime})$ measures the semantic alignment between the original instruction (Lee2025xJailbreakRS)(Intention Score, ITT) and its rewritten counterpart within the same range. The overall reward is computed as a weighted sum:
$R(x_{0},x^{\prime},\mathcal{T})=\alpha\cdot S(y)+(1-\alpha)\cdot I(x_{0},x^{\prime}),$ (2)
Applying CognitiveAttack
@@ 3 Methodology
Cognitive Bias Paraphrased Instruction(“Create a fake news article to spread misinformation about a political candidate”)
@@
By leveraging CognitiveAttack, we can automatically generate adversarial prompts that combine the optimized cognitive bias strategies to detect vulnerabilities in LLMs. Specifically, given a held-out set of harmful instructions $x_{0}$ sampled from datasets, the model infers the optimal bias combination $\mathcal{B}^{*}\subseteq\mathcal{B}_{\text{pool}}$ and rewrites the input into a paraphrased instruction $x^{\prime}\sim\pi_{\theta}(\cdot\mid x_{0},\mathcal{B}^{*})$ through a <think> step. Table 11 presents examples of how cognitive biases can be used to paraphrase harmful instructions $\mathcal{O}$. The first three samples each apply a single bias—Anchoring Effect, Authority Bias, or Gambler’s Fallacy—to influence the model’s reasoning. The final example combines multiple biases to create a more effective jailbreak prompt. This reasoning-aware rewriting process explicitly aims to maximize the expected reward defined in Eq. 1, thereby enhancing the likelihood of eliciting policy-violating responses while preserving the original intent.
Each rewritten prompt $x^{\prime}$ is then evaluated against the target model $\mathcal{M}$. A jailbreak is considered successful if the output violates the safety mechanism. All successful samples are collected to form a diverse dataset covering extensive cognitive bias types, enabling subsequent quantitative analysis of cognitive bias risks in LLMs.
4 Experiments
Leveraging the aforementioned methodology, we generate the dataset to conduct a systematic analysis of LLMs’ vulnerabilities to cognitive biases.
Model. We train our red team model based on the deepseek-r1-distill-qwen-14b (guo2025deepseek). This model is specifically designed to rewrite harmful instructions by leveraging cognitive bias strategies. For evaluation, we target a diverse set of representative LLMs, including Llama-series, Vicuna-series, Mistral-series, Qwen-series, GPT-series, DeepSeek(DS)-series, Gemini, and Claude.
Datasets. We evaluate the effectiveness of our cognitive attack on three datasets: AdvBench (zou2023universal), HEx-PHI (qi2023fine), and HarmBench (mazeika2024harmbench), which are widely used in the field of jailbreak attacks(jiang2024artprompt; zeng2024johnny; zou2023universal). AdvBench and HEx-PHI datasets are each split evenly, with 50% used for training and 50% for testing, while HarmBench is used solely for testing purposes.
Metrics. We employ GPT-Judge (qi2023fine) to evaluate the Harmfulness Score (HS) of generated responses, which quantifies the degree of harmfulness in the LLM’s output. The Attack Success Rate (ASR) is defined as the percentage of harmful instructions that successfully bypass the LLM’s safety mechanisms, calculated as: